131-Iodine has a half life of 8.1 days. How much of a 1800 gram sample remains after 20 days?


1
Expert's answer
2021-05-07T05:25:14-0400

325.0787381


N(t)= N(o)(1/2)^t/t1/2

Where N(t) = Remaining mass

N(o) =Original mass

T =Time

T1/2 = Half life

N(t) = 1800(1/2)^20/8.1

= 1800(0.5)^20/8.1

=325.0787381 grams